OpenStudy (anonymous):

which of the following equations represents a line that is not parallel to the line 2x+5-4y=2y-4 a. -y=x-5 B.2y=-2x-12 C. x=y-8 D. 1-4y=4x-172

OpenStudy (dmb1215):

so 6y=2x-9

OpenStudy (dmb1215):

sor +1

OpenStudy (dmb1215):

wait 2x+9=6y

OpenStudy (dmb1215):

the slope is 2 so it would be B.

OpenStudy (anonymous):

It is d

OpenStudy (dmb1215):

How in order for it to be perpendicular it has to be the negative recipcol

OpenStudy (anonymous):

idk

OpenStudy (dmb1215):

well i dont know then

OpenStudy (anonymous):

yeah i thought so

OpenStudy (anonymous):

loser

OpenStudy (dmb1215):

ok ill find it

OpenStudy (anonymous):

find what

OpenStudy (dmb1215):

ok make sure the question is right because im finding that none of the 4 are parralell

OpenStudy (anonymous):

flutter u

OpenStudy (dmb1215):

well sorry i could not help

OpenStudy (texaschic101):

2x + 5 - 4y = 2y - 4 -- subtract 2y from both sides 2x + 5 - 4y - 2y = -4 -- subtract 5 from both sides 2x - 4y - 2y = -4 - 5 -- subtract 2x from both sides -4y - 2y = -2x - 4 - 5 -- combine like terms -6y = -2x - 9 -- divide by -6 y = 1/3x + 3/2 slope is 1/3. So you want a slope other then 1/3. (a) - y = x - 5 -- multiply by -1 y = -x + 5 -- slope is -1 (b) 2y = -2x - 12 -- divide by 2 y = -x - 6 -- slope is -1 (c) x = y - 8 x + 8 = y or y = x + 8 -- slope is 1 (d) 1 - 4y = 4x - 172 -- subtract 1 from both sides -4y = 4x - 172 - 1 -4y = 4x - 173 -- divide by -4 y = -x + 173/4 --- slope is -1 well...here is the deal, none of your answer choices is parallel to the given line. And none of your answer choices are perpendicular to the given line.. However, the only one that does not match is answer choice C. It has a positive 1 slope and the other answer choices have a negative 1 slope.

OpenStudy (dmb1215):

i tried to say that but i was wrong i guess

OpenStudy (texaschic101):

if you tried to say that, then you are not wrong. The answer choices do not match the question. Let me rephrase that, all of the answer choices are correct. None of them are parallel to the given line

OpenStudy (dmb1215):

i told her none of the lines were parrelell to it

OpenStudy (texaschic101):

and you were correct :)

OpenStudy (dmb1215):

Well she/he thought i was stupid

OpenStudy (texaschic101):

you obviously are not. Some people are just not very grateful for what people on OS try to do for them.

OpenStudy (dmb1215):

I know thanks for the medal

OpenStudy (texaschic101):

no problem....oh, thanks for the medal, as well :)

OpenStudy (dmb1215):

No problem